OpenStudy (vishweshshrimali5):

A little, cute and nice question. Give it a try. Number theory problem.

OpenStudy (vishweshshrimali5):

Prove that \(\large{2^p + 3^p}\) is not a perfect power if p is a prime number.

mathslover (mathslover):

Use Mathematical Induction.. ?

OpenStudy (vishweshshrimali5):

@ganeshie8 , @mathslover , @Miracrown

OpenStudy (vishweshshrimali5):

Nope you can't use mathematical induction here. Because you don't have a series or sequence of numbers which are prime numbers.

OpenStudy (vishweshshrimali5):

^^ The very basic criterion required to apply PMI.

OpenStudy (ikram002p):

u mean u cant write \(2^p+3^p=(2+3)^n\) :D

mathslover (mathslover):

hmm, write 3 as 2 + 1...

OpenStudy (marissalovescats):

"A little, cute and nice question" >cute Lol

OpenStudy (vishweshshrimali5):

Yeah @ikram002p is right

OpenStudy (ikram002p):

binomial thm ^

mathslover (mathslover):

You get : \(2^ p + (2+1)^p \)

OpenStudy (vishweshshrimali5):

@marissalovescats : Of course its cute. See, how innocent is this question. It only has a 2, 3 and n. No, x or y etc.

mathslover (mathslover):

Use binomial theorem now.. yeah ikram, u r right.

OpenStudy (marissalovescats):

:P

OpenStudy (vishweshshrimali5):

Well guys go on give it a try with binomial theorem. Its just that I am not going to come to your rescue. :P

OpenStudy (anonymous):

Isn't that a theorem , it says that no power of a prime is a perfect number

OpenStudy (vishweshshrimali5):

@marissalovescats : Have you ever seen a more cuter problem ? :P

OpenStudy (anonymous):

Well, then i know to prove that

OpenStudy (vishweshshrimali5):

Nope: 2^4 + 3^2 = 25 = 5^2 Theorem fails

OpenStudy (marissalovescats):

If I ever see a math problem involving cats, I'll let you know :P

OpenStudy (vishweshshrimali5):

:P

OpenStudy (ikram002p):

meow @marissalovescats :P

OpenStudy (anonymous):

i saw a panda in p totaly cute <3

OpenStudy (marissalovescats):

I feel ya. It's always good to feel meow at 3:30am I wear a panda necklace daily. His name's Marty. He understands me.

OpenStudy (vishweshshrimali5):

Hey that was my next question @aajugdar . You spoiled it man. XD

OpenStudy (anonymous):

I meant no perfect number is a power of a prime.

OpenStudy (anonymous):

hehehe

OpenStudy (vishweshshrimali5):

Okay jokes apart friends. try to prove the theorem true for one exceptional prime number first.

mathslover (mathslover):

2^2 + 3^2 = 4 + 9 = 13..........! (Exceptional prime number)

OpenStudy (marissalovescats):

I'll never know :( "Three guests check into a hotel room. The clerk says the bill is $30, so each guest pays $10. Later the clerk realizes the bill should only be $25. To rectify this, he gives the bellhop $5 to return to the guests. On the way to the room, the bellhop realizes that he cannot divide the money equally. As the guests didn't know the total of the revised bill, the bellhop decides to just give each guest $1 and keep $2 as a tip for himself. Each guest got $1 back: so now each guest only paid $9; bringing the total paid to $27. The bellhop has $2. And $27 + $2 = $29 so, if the guests originally handed over $30. What happened to the remaining hour?"

OpenStudy (ikram002p):

okk so forget bi 2^p+3^p=(2+3)^n =5^n >.>

OpenStudy (ikram002p):

so if i proved that 2^p+3^p can devide any number not only 5 then done

OpenStudy (anonymous):

@marissalovescats the way of thinking is wrong in this one they took different perspective to trick you you will get it once you actualy try it by urself

OpenStudy (marissalovescats):

I've spend much time in my life trying to figure it out

OpenStudy (anonymous):

hehehe 27 is the charge they are paying extra 2 Rs where 25 is actualy charge :P so 27-25 = 2 $ which the guy took :D

OpenStudy (anonymous):

perspective :P

OpenStudy (anonymous):

you are considering 2 Rs for the 3 Rs he gave them back xD

ganeshie8 (ganeshie8):

@ikram002p try FLT

OpenStudy (ikram002p):

im wondering where is it , its abt n^p-1=1 mod p if gcd (p,n) =1 right?

ganeshie8 (ganeshie8):

its going to be an one line proof :P

OpenStudy (ikram002p):

haha ohk i got it nw 2^p+3^p mod p = 2+3 mod p then what ?

OpenStudy (anonymous):

@marissalovescats simple way to know would be 30 rs is what they paid he gave them back 3 rs 30-3 =27 n he took 2 Rs for himself 27-2 = 25 but for them they paid 30 rs for which they got 3 Rs back so it will be 27+3 and not 27+2 tricky tricky

ganeshie8 (ganeshie8):

\[\large 2^p \equiv 2 \mod p \implies 2^p = pu+ 2\] \[\large 3^p \equiv 3\mod p \implies 3^p = pv +3\]

ganeshie8 (ganeshie8):

set it equal to a power and show that no solutions exist

OpenStudy (ikram002p):

continue plz

OpenStudy (marissalovescats):

Dang that reallly messes with my mind

OpenStudy (anonymous):

hehe

OpenStudy (gudden):

Cutest OS question .. .:)

OpenStudy (ikram002p):

\(2^p+3^p=5^n\) take mod p \((2^p+3^p) \mod p=5^n \mod p\) \((2+3) \mod p=5^n \mod p\) \(5 \mod p=5^n \mod p\) so n=1 mmm thats it ?

ganeshie8 (ganeshie8):

Looks perfect !

OpenStudy (ikram002p):

that remind me with fermat last thm \(a^n+b^n=c^2\) has no solution for n >2 but i guess its not proven yet ( or idk)

ganeshie8 (ganeshie8):

lol you're missing the fun all these days ! google up wiley

ganeshie8 (ganeshie8):

*wiles proof

OpenStudy (ikram002p):

wait ill check

ganeshie8 (ganeshie8):

watch this sequence when free :) https://www.youtube.com/watch?v=PPOhFok6H-Q

OpenStudy (ikram002p):

lol ok ok ive seen it before haha

OpenStudy (marissalovescats):

Still qtest OS qeustion 2k14

OpenStudy (ikram002p):

ok @vishweshshrimali5 what is ur proof ?

OpenStudy (gudden):

Everybody...! Now help me with my question http://openstudy.com/study#/updates/539d53a9e4b0206eed099306

OpenStudy (vishweshshrimali5):

Ok friends here is my solution : (1) As @mathslover did, I am going to prove the given statement true for p = 2 (only even prime number and hence an exceptional one) \(\large{2^2 + 3^2 = 13 \ne \text{perfect power}}\) (2) Now all prime numbers more than 2 are odd. Now by factor theorem, \(\large{(x+a)|(x^p + a^p)}\) Which means that (x+a) is a factor of (x^p + a^p) when p is odd. So, \(\large{5|(2^p + 3^p)}\) Now, for proving that this is not a perfect number, I am going to prove that it is NOT divisible by 5^2. Now, \(\large{(x^p + 3^p) = (x+3)(x^{p-1} - 3x^{p-2} + ... + (-3)^{p-1})}\) Now, when x = -3 then, \(\large{(x^{p-1} - 3x^{p-2} + ... + (-3)^{p-1})}\) \(\large{ = (-3)^{p-1} - 3 * (-3)^{p-2} + ... + (-3)^{p-1}}\) \(\large{ = p*3^{p-1}}\) (Since, p is odd, so, p-1 is even, p-2 is odd and so on) \(\large{ \ne 0}\) So, clearly, (x+3) is not a factor of : \(\large{(x^{p-1} - 3x^{p-2} + ... + (-3)^{p-1})}\) (factor theorem) Thus, \(\large{(x+3)^2}\) is not a factor of \(\large{x^p + 3^p}\). Now, by replacing x by 2, I get that, \(\large{(2+3)^2 = 5^2}\) is not a factor of \(\large{2^p + 3^p}\). Thus, \(\large{2^p + 3^p}\) is not a perfect power for prime p.

OpenStudy (vishweshshrimali5):

@ganeshie8 @hartnn @ikram002p @mathslover @Miracrown

OpenStudy (vishweshshrimali5):

@ikram002p what do you think ?

OpenStudy (vishweshshrimali5):

I think this can be generalised for any x and a provided that a or x are not divisible by each other. Thus, \(\large{x^p + a^p}\) is not a perfect power if p is a prime number and (x,a) = 1.

OpenStudy (ikram002p):

great ^_^ nice work

OpenStudy (vishweshshrimali5):

Thanks @ikram002p

OpenStudy (ikram002p):

np , ur wlc :)

OpenStudy (vishweshshrimali5):

:)

ganeshie8 (ganeshie8):

very clever application of factor theorem :)

OpenStudy (vishweshshrimali5):

thanks @ganeshie8